PROOF OF A DETERMINANT EVALUATION CONJECTURED BY BOMBIERI, HUNT AND VAN DER POORTEN

CHRISTIAN KRATTENTHALER AND DORON ZEILBERGER

Abstract. A determinant evaluation is proven, a special case of which establishes a conjecture of Bombieri, Hunt, and van der Poorten (Experimental Math. 4 (1995), 87–96) that arose in the study of Thue’s method of approximating algebraic numbers.